### v3

@@ 1 Introduction
For $k=3$, deciding whether a graph contains a triangle and finding one if it does, or counting all triangles in a graph, can be done in $O(n^{\omega})$ time by the algorithm of Itai and Rodeh [16], where $\omega$ is the exponent of matrix multiplication [1, 7], and $\omega<2.372$ is the best known bound [10, 38]. The algorithm compares $M$ and $N=M^{2}$, where $M$ is the graph adjacency matrix; if there is a pair of indexes $i,j$ such that $M[i,j]=N[i,j]=1$, then there is a triangle based on edge $ij$. An equivalent characterization is: $G$ contains a triangle iff $M^{3}$ has a non-zero entry on its main diagonal. See [27, Ch. 10] for a short exposition of this elegant method. Alternatively, this task can be done in $O(m^{2\omega/(\omega+1)})=O(m^{1.407})$ time by the algorithm of Alon, Yuster, and Zwick [3], which deals with vertices of high and low degree separately. Itai and Rodeh [16] and also Papadimitriou and Yannakakis [32] as well as Chiba and Nishizeki [6] showed that triangles in planar graphs can be found in $O(n)$ time.
@@
Chiba and Nishizeki [6] further refined the time complexity in terms of graph arboricity, which is the minimum number of edge-disjoint forests into which its edges can be partitioned. Their algorithm lists all triangles in a graph in $O(m\alpha)$ time, where $\alpha$ is the arboricity of the graph. They also showed that $\alpha=O(\sqrt{m})$; consequently, the running time is $O(m^{3/2})$. Since there are graphs $G$ with $\alpha(G)=\Theta(m^{1/2})$, this does not improve the worst-case dependence on $m$. Moreover, since there are graphs with $\Theta(m^{3/2})$ triangles, see, e.g., [33], this dependence cannot be improved. For every fixed $\ell\geq 3$, the same authors gave an algorithm for listing all copies of $K_{\ell}$ in $O(\alpha^{\ell-2}\cdot m)=O(m^{\ell/2})$ time. If $G$ is planar, $\alpha(G)\leq 3$ (see [15, p. 124]), so their algorithm runs in $O(m)=O(n)$ (i.e., linear) time on planar graphs.
@@
It should be noted that the asymptotically fastest algorithms for matrix multiplication have mainly a theoretical importance and are otherwise largely impractical [26, Ch. 10.2.4]. As such, any algorithm that employs fast matrix multiplication (FMM) inherits this undesirable feature. Whereas the notion of “combinatorial” algorithm does not have a formal or precise definition, such an algorithm is usually simpler to implement and practically efficient, see [37]. Combinatorial algorithms are sometimes theoretically less efficient than non-combinatorial ones but practically more efficient since they tend to be simpler to implement.
@@
For a graph $G$, its arboricity $\alpha(G)$ is the minimum number of edge-disjoint forests into which $G$ can be decomposed [6]. For instance, it is known and easy to show that $\alpha(K_{n})=\lceil n/2\rceil$. A characterization of arboricity is provided by the following classic result [28, 29, 36]; see also [9, p. 60].
@@
(Nash-Williams 1964; Tutte 1961) A multigraph $G=(V,E)$ can be partitioned into at most $k$ forests if and only if every set $U\subseteq V$ induces at most $k(|U|-1)$ edges.
@@ 3 Delete the tree edges from $G$ and go to Step 1
The second is a randomized algorithm that checks whether there is an edge contained in a triangle. It runs in $O(mn)$ worst-case time and $O(n^{5/3})$ expected time. The third algorithm relies on Boolean matrix multiplication and runs in $O(n^{\omega})$ time. (Boolean matrix multiplication can be reduced to standard matrix multiplication; see [26, Ch. 10.2.4] for details.)
@@ 1 Sort vertices such that $\deg(v_{1})\geq\deg(v_{2})\geq\ldots\deg(v_{n})$
6                if $w$ is marked then output triangle $uvw$
@@ 2 A Simple Hybrid Algorithm for Listing All Triangles
In this section we present a new algorithm for listing all triangles. While its general idea is not new, the specific hybrid data structure therein does not appear to have been previously considered. Using both an adjacency list representation and an adjacency matrix representation of the graph yields a very time-efficient neighborhood merge (intersection) procedure. Let $V=\{1,2,\ldots,n\}$ and let $M$ be the adjacency matrix of $G$. A triangle $ijk$ with $i<j<k$ is reported when edge $ij$ is processed, and so each triangle is reported exactly once. For each edge $ij$, the algorithm scans the adjacency list of the endpoint of lower degree (among $i$ and $j$) and for each neighbor it checks for the needed entry in the adjacency matrix $M$ corresponding to the third triangle edge. In the pseudocode below, $ADJ(v)$ denotes the adjacency list of vertex $v$.
Hybrid$(G)$
@@
5           if $j<k$ and $M(y,k)=1$ then report triangle $ijk$
@@
We subsequently show that the algorithm runs in time $O(m\alpha)$. Whereas the space used is quadratic, the hybrid algorithm appears to win by running time and simplicity. In particular, no additional data structures nor hashing are used, no sorting (by degree or otherwise) and no doubly linked lists are needed, etc. Similar algorithms that use hashing in order to run in linear space are $6$ to $8$ times slower than their “unoptimized” counterparts on most graphs with about $5\times 10^{6}$ vertices in the study in [34, 35]. “The two algorithms edge-iterator-hashed and forward-hashed, which use hashed data structures for every node, perform badly” on such graphs, the authors write [35, p. 8].
@@ 2.1 The Analysis
There are at most $f(e)$ triangles based on edge $e=ij$ and overall at most $F(G)$ triangles in $G$. The runtime of the algorithm is proportional to this quantity; the space used is quadratic. A short and elegant decomposition argument by Chiba and Nishizeki [6, Lemma 2] shows that
@@
Each edge of type 2 connects two nodes of degree $>2\sqrt{m}$, and there are at most $2m/(2\sqrt{m})=\sqrt{m}$ such nodes. The degree of each of them thus contributes to $F(G)$ at most $\sqrt{m}$ times and the sum of all degrees of such nodes is at most $2m$. It follows that the total contribution from edges of type 2 is at most $2m^{3/2}$.
@@
Lemma 1 immediately gives an upper bound of $4m^{3/2}$ on the number of triangles in a graph. It is in fact known [33] that this number is at most $(2^{1/2}/3)\,m^{3/2}$ (and this bound is sharp). More generally, for every fixed $\ell\geq 3$, the number of copies of $K_{\ell}$ is $O(\alpha^{\ell-2}\cdot m)=O(m^{\ell/2})$; see [6, Thm. 3].
@@ 3 Constructions
Let $3\leq x\leq n$ be the unique integer such that (${x\choose 2}\leq m<{x+1\choose 2}$. Note that $x=\Theta(\sqrt{m})$. Let $G=(V,E)$ where $V=V_{1}\cup V_{2}$, $|V_{1}|=x$, and $|V_{2}|=n-x$. Set $V_{1}$ induces a complete subgraph and set $V_{2}$ induces an independent set in $G$. Add $m-{x\choose 2}$ edges to $G$ arbitrarily so that $G$ has $m$ edges. Let $T$ be the set of triangles in $G$. We have
@@
We may assume that $b$ is even and $b/2\ |\ n$. Suppose that $n=(2k+1)b/2$. Let $G=(V,E)$, where $V=V_{1}\cup V_{2}$, and $|V_{1}|=kb$, $|V_{2}|=b/2$. Set $V_{1}$ consists of $k$ complete subgraphs of order $b$ whereas $V_{2}$ is an independent set of size $b/2$ in $G$. All edges in $V_{1}\times V_{2}$ are present in $G$. See Fig. 1.
@@ 4 Finding Small Complete Subgraphs Efficiently
The algorithm solves the detection (or counting) problem for a complete subgraph $K_{\ell}$. Let $T(n,m,\ell)$ denote the running time of the algorithm for a graph with $n$ vertices and $m$ edges. Write $\ell=\ell_{1}+\ell_{2}$, for a suitable choice $\ell_{1},\ell_{2}\geq 2$. At the beginning, we run the algorithm of Chiba and Nishizeki to form a list of subgraphs isomorphic to $K_{\ell_{1}}$. Then, for each subgraph $G_{1}=(V_{1},E_{1})$ on the list: (i) we construct the subgraph $G_{2}$ of $G$ induced by vertices in $V\setminus V_{1}$ that are adjacent to all vertices in $V_{1}$; and (ii) we count the subgraphs isomorphic to $K_{\ell_{2}}$ in $G_{2}$ (or find one such subgraph); this is a (possibly recursive) call of the same procedure on a smaller instance. In other words, we count the number of extensions of the subgraph isomorphic to $K_{\ell_{1}}$ to a $K_{\ell}$. Another formulation of this method can be found in [21].
There are $O(\alpha^{\ell_{1}-2}\cdot m)$ copies of $K_{\ell_{1}}$ and they can be found in $O(\alpha^{\ell_{1}-2}\cdot m)$ time. For each fixed copy of $K_{\ell_{1}}$, the time spent in $G_{2}$ is at most $T(n,m,\ell_{2})$, and so the overall time satisfies the recurrence
$T(n,m,\ell)=O(\alpha^{\ell_{1}-2}\cdot m\cdot T(n,m,\ell_{2})).$
Each copy of $K_{\ell}$ is generated exactly ${\ell\choose\ell_{1}}$ times and so the total count needs to be divided by this number in the end.
@@
Here we focus on $k=3$. For convenience, define the following integer functions
@@
With this notation, the algorithm runs in $O(n^{\gamma(\ell)})$ time. Two decades later, Eisenbrand and Grandoni [12] refined the triangle method by using fast algorithms for rectangular matrix multiplication instead of those for square matrix multiplication. It partitions the graph into three parts of roughly the same size: $\ell_{1}=\lfloor\ell/3\rfloor$, $\ell_{2}=\lceil(\ell-1)/3\rceil$, $\ell_{3}=\lceil\ell/3\rceil$. The refined algorithm runs in time $O(n^{\beta(\ell)})$ time. If the rectangular matrix multiplication is carried out via the straightforward partition into square blocks and fast square matrix multiplication (see, e.g., [8, Exercise 4.2-6]), one recovers the time complexity of the algorithm of Nešetřil and Poljak; that is, $\beta(\ell)\leq\gamma(\ell)$, see [12] for details. Eisenbrand and Grandoni [12] showed that the above inequality is strict for a certain range: if $\ell\pmod{3}=1$ and $\ell\leq 16$, or $\ell\pmod{3}=2$. In summary, their algorithm is faster than that of Nešetřil and Poljak in these cases. We subsequently refer to their method as the refined triangle method. Another extension of the triangle method is considered in [21].
@@
    $\ell=3j+1$, $j\geq 2$. The refined triangle method [12] with $\ell_{1}=j$, $\ell_{2}=j+1$, $\ell_{3}=j$, leads to rectangular matrix multiplication $[O(\alpha^{j-2}m)\times O(\alpha^{j-1}m)]\cdot[O(\alpha^{j-1}m)\times O(\alpha^{j-2}m)]$. Its complexity is at most $O(\alpha)$ times that of the square matrix multiplication with dimension $\alpha^{j-2}m$, the latter of which is $O(\alpha^{(j-2)\omega}\cdot m^{\omega})$. It follows that
@@
    $\ell=3j+2$, $j\geq 2$. The refined triangle method [12] with $\ell_{1}=j+1$, $\ell_{2}=j$, $\ell_{3}=j+1$, leads to rectangular matrix multiplication $[O(\alpha^{j-1}m)\times O(\alpha^{j-2}m)]\cdot[O(\alpha^{j-2}m)\times O(\alpha^{j-1}m)]$. Its complexity is at most $O(\alpha^{2})$ times that of the square matrix multiplication with dimension $\alpha^{j-2}m$, the latter of which is $O(\alpha^{(j-2)\omega}\cdot m^{\omega})$. It follows that
@@
$3$ $O(\alpha\cdot m)$ [6], $O(n^{2.372})$ [16], $O(m^{1.407})$ [3]
$4$ $O(\alpha^{2}\cdot m)$ [6], $O(n^{3.334})$ [12], $O(m^{1.686})$ [18] $O(n^{3.251})$
$5$ $O(\alpha^{3}\cdot m)$ [6], $O(n^{4.220})$, $O(m^{2.147})$ [12] $O(n^{4.086})$
$6$ $O(\alpha^{4}\cdot m)$ [6], $O(n^{4.751})$ [12], $O(m^{2.372})$ [12]
$7$ $O(\alpha^{5}\cdot m)$ [6], $O(n^{5.714})$ [12], $O(m^{2.857})$ [12] $O(m^{2.795})$, $O(\alpha\cdot m^{2.372})$
$8$ $O(\alpha^{6}\cdot m)$ [6], $O(m^{3.372})$ [12, 18] $O(m^{3.251})$, $O(\alpha^{3.372}\cdot m^{1.686})$
$9$ $O(\alpha^{7}\cdot m)$ [6], $O(m^{3.558})$ [12, 18] $O(\alpha^{2.372}\cdot m^{2.372})$
@@
$14$ $O(\alpha^{11}\cdot m)$ [6], $O(m^{5.553})$ [12, 18] $O(\alpha^{6.744}\cdot m^{2.372})$
@@
    $\ell=4$. The refined triangle method [12] with $\ell_{1}=1$, $\ell_{2}=1$, $\ell_{3}=2$, leads to rectangular matrix multiplication $[n\times n]\cdot[n\times m]$ or $[n\times n]\cdot[n\times n^{2}]$ in the worst case. Since $\omega(1,1,2)=\omega(1,2,1)\leq 3.251$, it follows that $T(n,m,4)=O(n^{3.251})$. (See also [22, Table 3]; this entry has been slightly improved in [38].) In particular, this gives a positive answer to a question of Alon, Yuster, and Zwick [3], who asked for an improvement of their $O(n^{\omega+1})=O(n^{3.372})$ upper bound for $K_{4}$ detection. In terms of $m$, by [12, Table 1], $T(n,m,4)=O(m^{1.682})$, but this entry appears unjustified.
@@
    $\ell=5$. The refined triangle method [12] with $\ell_{1}=2$, $\ell_{2}=1$, $\ell_{3}=2$, leads to rectangular matrix multiplication $[m\times n]\cdot[n\times m]$ or $[n^{2}\times n]\cdot[n\times n^{2}]$ in the worst case. According to [38, Table 1], $\omega(2,1,2)=2\omega(1,0.5,1)\leq 2\cdot 2.043=4.086$, hence $T(n,m,5)=O(n^{4.086})$.
    On a side note, the extension method [12] with $\ell_{1}=3,\ell_{2}=2$, yields $T(n,m,5)=O(\alpha m\cdot m)=O(\alpha\cdot m^{2})$. Observe however, that $O(\alpha^{3}\cdot m)=O(\alpha\cdot m^{2})$, so the above upper bound does not beat that of the algorithm of Chiba and Nishizeki.
@@
    $\ell=7$. The refined triangle method [12] with $\ell_{1}=2$, $\ell_{2}=3$, $\ell_{3}=2$, leads to rectangular matrix multiplication $[m\times\alpha m]\cdot[\alpha m\times m]$ or $[m\times m^{3/2}]\cdot[m^{3/2}\times m]$ in the worst case. Since $\omega(1,1.5,1)\leq 2.795$, it follows that $T(n,m,7)=O(m^{2.795})$.
@@
    $\ell=8$. The refined triangle method [12] with $\ell_{1}=\ell_{2}=\ell_{3}=\ell_{4}=2$, leads to counting $K_{4}$’s in a graph with $m$ vertices. Since $\omega(1,1,2)\leq 3.251$, we have $T(n,m,8)=O(m^{3.251})$.
@@
    $\ell=10$. By [12, Thm. 2], it takes $O(m^{\beta(10)/2})$, where $\beta(10)=\omega(3,3,4)$. Since according to [22, Table 3], $\omega(1,1.3,1)\leq 2.6217$ and $\omega(1,1.4,1)\leq 2.7084$, by the convexity of $\omega(1,x,1)$ (see [24]), we have $\omega(1,4/3,1)=\omega(1,1.3*2/3+1.4/3,1)\leq 2\omega(1,1.3,1)/3+\omega(1,1.4,1)/3\leq 7.9518/3$. Then $\omega(3,3,4)=3\omega(1,4/3,1)\leq 7.952$, thus $T(n,m,10)=O(m^{\beta(10)/2})=O(m^{3.976})$. Example: for graphs with $\alpha=O(m^{3/8})$, $K_{10}$’s can be counted/detected in $O(m^{3.636})$ time, by the last entry in the right column, whereas the other entries give $O(m^{3.976})$ time or more.
@@
    $\ell=11$. By [12, Thm. 2], it takes $O(m^{\beta(11)/2})$, where $\beta(11)=\omega(3,4,4)$. By [38, Table 1], we have $\omega(1,0.75,1)\leq 2.1863$, which yields $\omega(3,4,4)=4\omega(1,0.75,1)\leq 4\times 2.1863\leq 8.746$. Consequently, $T(n,m,11)=O(m^{\beta(11)/2})=O(m^{4.373})$.

### v4

@@ 1 Introduction
For $k=3$, deciding whether a graph contains a triangle and finding one if it does, or counting all triangles in a graph, can be done in $O(n^{\omega})$ time by the algorithm of Itai and Rodeh [17], where $\omega$ is the exponent of matrix multiplication [2, 8], and $\omega<2.372$ is the best known bound [11, 39]. The algorithm compares $M$ and $N=M^{2}$, where $M$ is the graph adjacency matrix; if there is a pair of indexes $i,j$ such that $M[i,j]=N[i,j]=1$, then there is a triangle based on edge $ij$. An equivalent characterization is: $G$ contains a triangle iff $M^{3}$ has a non-zero entry on its main diagonal. See [28, Ch. 10] for a short exposition of this elegant method. Alternatively, this task can be done in $O(m^{2\omega/(\omega+1)})=O(m^{1.407})$ time by the algorithm of Alon, Yuster, and Zwick [4], which deals with vertices of high and low degree separately. Itai and Rodeh [17] and also Papadimitriou and Yannakakis [33] as well as Chiba and Nishizeki [7] showed that triangles in planar graphs can be found in $O(n)$ time.
@@
Chiba and Nishizeki [7] further refined the time complexity in terms of graph arboricity, which is the minimum number of edge-disjoint forests into which its edges can be partitioned. Their algorithm lists all triangles in a graph in $O(m\alpha)$ time, where $\alpha$ is the arboricity of the graph. They also showed that $\alpha=O(\sqrt{m})$; consequently, the running time is $O(m^{3/2})$. Since there are graphs $G$ with $\alpha(G)=\Theta(m^{1/2})$, this does not improve the worst-case dependence on $m$. Moreover, since there are graphs with $\Theta(m^{3/2})$ triangles, see, e.g., [34], this dependence cannot be improved. For every fixed $\ell\geq 3$, the same authors gave an algorithm for listing all copies of $K_{\ell}$ in $O(\alpha^{\ell-2}\cdot m)=O(m^{\ell/2})$ time. If $G$ is planar, $\alpha(G)\leq 3$ (see [16, p. 124]), so their algorithm runs in $O(m)=O(n)$ (i.e., linear) time on planar graphs.
@@
It should be noted that the asymptotically fastest algorithms for matrix multiplication have mainly a theoretical importance and are otherwise largely impractical [27, Ch. 10.2.4]. As such, any algorithm that employs fast matrix multiplication (FMM) inherits this undesirable feature. Whereas the notion of “combinatorial” algorithm does not have a formal or precise definition, such an algorithm is usually simpler to implement and practically efficient, see [38]. Combinatorial algorithms are sometimes theoretically less efficient than non-combinatorial ones but practically more efficient since they tend to be simpler to implement.
@@
For a graph $G$, its arboricity $\alpha(G)$ is the minimum number of edge-disjoint forests into which $G$ can be decomposed [7]. For instance, it is known and easy to show that $\alpha(K_{n})=\lceil n/2\rceil$. A characterization of arboricity is provided by the following classic result [29, 30, 37]; see also [10, p. 60].
@@
(Nash-Williams 1964; Tutte 1961) A multigraph $G=(V,E)$ can be partitioned into at most $k$ forests if and only if every set $U\subseteq V$ induces at most k(|U|1)k(|U|-1) edges.
@@ 3 Delete the tree edges from $G$ and go to Step 1
The second is a randomized algorithm that checks whether there is an edge contained in a triangle. It runs in $O(mn)$ worst-case time and $O(n^{5/3})$ expected time. The third algorithm relies on Boolean matrix multiplication and runs in $O(n^{\omega})$ time. (Boolean matrix multiplication can be reduced to standard matrix multiplication; see [27, Ch. 10.2.4] for details.)
@@ 1 Sort vertices such that $\deg(v_{1})\geq\deg(v_{2})\geq\ldots\deg(v_{n})$
6       if $w$ is marked then output triangle uvwuvw
@@ 2 A Simple Hybrid Algorithm for Listing All Triangles
In this section we present a new algorithm for listing all triangles. While its general idea is not new, the specific hybrid data structure therein does not appear to have been previously considered. Using both an adjacency list representation and an adjacency matrix representation of the graph yields a very time-efficient neighborhood merge (intersection) procedure. Let $V=\{1,2,\ldots,n\}$ and let $M$ be the adjacency matrix of $G$. A triangle ijkijk with $i<j<k$ is reported when edge $ij$ is processed, and so each triangle is reported exactly once. For each edge $ij$, the algorithm scans the adjacency list of the endpoint of lower degree (among $i$ and $j$) and for each neighbor it checks for the needed entry in the adjacency matrix $M$ corresponding to the third triangle edge. In the pseudocode below, $ADJ(v)$ denotes the adjacency list of vertex $v$.
Hybrid(G)(G)
@@
5     if $j<k$ and $M(y,k)=1$ then report triangle ijkijk
@@
We subsequently show that the algorithm runs in time $O(m\alpha)$. Whereas the space used is quadratic, the hybrid algorithm appears to win by running time and simplicity. In particular, no additional data structures nor hashing are used, no sorting (by degree or otherwise) and no doubly linked lists are needed, etc. Similar algorithms that use hashing in order to run in linear space are $6$ to $8$ times slower than their “unoptimized” counterparts on most graphs with about $5\times 10^{6}$ vertices in the study in [35, 36]. “The two algorithms edge-iterator-hashed and forward-hashed, which use hashed data structures for every node, perform badly” on such graphs, the authors write [36, p. 8].
@@ 2.1 The Analysis
There are at most $f(e)$ triangles based on edge $e=ij$ and overall at most $F(G)$ triangles in $G$. The runtime of the algorithm is proportional to this quantity; the space used is quadratic. A short and elegant decomposition argument by Chiba and Nishizeki [7, Lemma 2] shows that
@@
Each edge of type 2 connects two nodes of degree $2\sqrt{m}$, and there are at most $2m/(2\sqrt{m})=\sqrt{m}$ such nodes. The degree of each of them thus contributes to $F(G)$ at most $\sqrt{m}$ times and the sum of all degrees of such nodes is at most $2m$. It follows that the total contribution from edges of type 2 is at most $2m^{3/2}$.
@@
Lemma 1 immediately gives an upper bound of $4m^{3/2}$ on the number of triangles in a graph. It is in fact known [34] that this number is at most $(2^{1/2}/3)\,m^{3/2}$ (and this bound is sharp). More generally, for every fixed $\ell\geq 3$, the number of copies of $K_{\ell}$ is $O(\alpha^{\ell-2}\cdot m)=O(m^{\ell/2})$; see [7, Thm. 3].
@@ 3 Constructions
Let $3\leq x\leq n$ be the unique integer such that ${x\choose 2}\leq m<{x+1\choose 2}$. Note that $x=\Theta(\sqrt{m})$. Let $G=(V,E)$ where $V=V_{1}\cup V_{2}$, $|V_{1}|=x$, and $|V_{2}|=n-x$. Set $V_{1}$ induces a complete subgraph and set $V_{2}$ induces an independent set in $G$. Add $m-{x\choose 2}$ edges to $G$ arbitrarily so that $G$ has $m$ edges. Let $T$ be the set of triangles in $G$. We have
@@
We may assume that $b$ is even and b/2|nb/2\ |\ n. Suppose that $n=(2k+1)b/2$. Let $G=(V,E)$, where $V=V_{1}\cup V_{2}$, and $|V_{1}|=kb$, $|V_{2}|=b/2$. Set $V_{1}$ consists of $k$ complete subgraphs of order $b$ whereas $V_{2}$ is an independent set of size $b/2$ in $G$. All edges in $V_{1}\times V_{2}$ are present in $G$. See Fig. 1.
@@ 4 Finding Small Complete Subgraphs Efficiently
The algorithm solves the detection (or counting) problem for a complete subgraph $K_{\ell}$. Let $T(n,m,\ell)$ denote the running time of the algorithm for a graph with $n$ vertices and $m$ edges. Write $\ell=j_{1}+j_{2}$, for a suitable choice $j_{1},j_{2}\geq 2$. At the beginning, we run the algorithm of Chiba and Nishizeki to form a list of subgraphs isomorphic to $K_{j_{1}}$. Then, for each subgraph $G_{1}=(V_{1},E_{1})$ on the list: (i) we construct the subgraph $G_{2}$ of $G$ induced by vertices in $V\setminus V_{1}$ that are adjacent to all vertices in $V_{1}$; and (ii) we count the subgraphs isomorphic to $K_{j_{2}}$ in $G_{2}$ (or find one such subgraph); this is a (possibly recursive) call of the same procedure on a smaller instance. In other words, for each subgraph isomorphic to $K_{j_{1}}$, we count the number of extensions to $K_{\ell}$. Another formulation of this method can be found in [22].
There are $O(\alpha^{j_{1}-2}\cdot m)$ copies of $K_{j_{1}}$ and they can be found in $O(\alpha^{j_{1}-2}\cdot m)$ time. For each fixed copy of $K_{j_{1}}$, the time spent in $G_{2}$ is at most $T(n,m,j_{2})$, and so the overall time satisfies the recurrence
$T(n,m,\ell)=O(\alpha^{j_{1}-2}\cdot m\cdot T(n,m,j_{2})).$
Each copy of $K_{\ell}$ is generated exactly $\ell\choose j_{1}$} times and so the total count needs to be divided by this number in the end.
@@
Here we focus on $k=3$. For convenience, define the following functions on integer inputs
@@
With this notation, the algorithm runs in $O(n^{\gamma(\ell)})$ time. Two decades later, Eisenbrand and Grandoni [13] refined the triangle method by using fast algorithms for rectangular matrix multiplication instead of those for square matrix multiplication. It partitions the graph into three parts of roughly the same size: $\ell_{1}=\lfloor\ell/3\rfloor$, $\ell_{2}=\lceil(\ell-1)/3\rceil$, $\ell_{3}=\lceil\ell/3\rceil$. The refined algorithm runs in time $O(n^{\beta(\ell)})$ time. If the rectangular matrix multiplication is carried out via the straightforward partition into square blocks and fast square matrix multiplication (see, e.g., [9, Exercise 4.2-6]), one recovers the time complexity of the algorithm of Nešetřil and Poljak; that is, $\beta(\ell)\leq\gamma(\ell)$, see [13] for details. Eisenbrand and Grandoni [13] showed that the above inequality is strict for a certain range: if $\ell\pmod{3}=1$ and $\ell\leq 16$, or $\ell\pmod{3}=2$. In summary, their algorithm is faster than that of Nešetřil and Poljak in these cases. We subsequently refer to their method as the refined triangle method. Another extension of the triangle method is considered in [22].
@@
    $\ell=3j+1$, $j\geq 2$. The refined triangle method [13] with $\ell_{1}=j$, $\ell_{2}=j$, $\ell_{3}=j+1$, leads to rectangular matrix multiplication $O(\alpha^{j-2}m)\times O(\alpha^{j-2}m)]\cdot[O(\alpha^{j-2}m)\times O(\alpha^{j-1}m)]$. Its complexity is at most $O(\alpha)$ times that of the square matrix multiplication with dimension $\alpha^{j-2}m$, the latter of which is $O(\alpha^{(j-2)\omega}\cdot m^{\omega})$. It follows that
@@
    $\ell=3j+2$, $j\geq 2$. The refined triangle method [13] with $\ell_{1}=j$, $\ell_{2}=j+1$, $\ell_{3}=j+1$, leads to rectangular matrix multiplication $O(\alpha^{j-2}m)\times O(\alpha^{j-1}m)]\cdot[O(\alpha^{j-1}m)\times O(\alpha^{j-1}m)]$. Its complexity is at most $O(\alpha^{2})$ times that of the square matrix multiplication with dimension $\alpha^{j-2}m$, the latter of which is $O(\alpha^{(j-2)\omega}\cdot m^{\omega})$. It follows that
@@
33 $O(\alpha\cdot m)$ [7], $O(n^{2.372})$ [17], $O(m^{1.407})$ [4]
44 $O(\alpha^{2}\cdot m)$ [7], $O(n^{3.334})$ [13], $O(m^{1.686})$ [19] $O(n^{3.251})$
55 $O(\alpha^{3}\cdot m)$ [7], $O(n^{4.220})$, $O(m^{2.147})$ [13] $O(n^{4.086})$
66 $O(\alpha^{4}\cdot m)$ [7], $O(n^{4.751})$ [13], $O(m^{2.372})$ [13]
77 $O(\alpha^{5}\cdot m)$ [7], $O(n^{5.714})$ [13], $O(m^{2.857})$ [13] $O(m^{2.795})$, $O(\alpha\cdot m^{2.372})$
88 $O(\alpha^{6}\cdot m)$ [7], $O(m^{3.372})$ [13, 19] $O(m^{3.199})$, $O(\alpha^{3.372}\cdot m^{1.686})$
99 $O(\alpha^{7}\cdot m)$ [7], $O(m^{3.558})$ [13, 19] $O(\alpha^{2.372}\cdot m^{2.372})$
@@
$14$ $O(\alpha^{11}\cdot m)$ [7], $O(m^{5.553})$ [13, 19] $O(m^{5.552})$, $O(\alpha^{6.744}\cdot m^{2.372})$
@@
    $\ell=4$. The refined triangle method [13] with $\ell_{1}=1$, $\ell_{2}=1$, $\ell_{3}=2$, leads to rectangular matrix multiplication $n\times n]\cdot[n\times m]$ or $n\times n]\cdot[n\times n^{2}]$ in the worst case. Since $\omega(1,1,2)=\omega(1,2,1)\leq 3.251$, it follows that $T(n,m,4)=O(n^{3.251})$. (See also [23, Table 3]; this entry has been slightly improved in [39].) In particular, this gives a positive answer to a question of Alon, Yuster, and Zwick [4], who asked for an improvement of their $O(n^{\omega+1})=O(n^{3.372})$ upper bound for $K_{4}$ detection. In terms of $m$, by [13, Table 1], $T(n,m,4)=O(m^{1.682})$, but this entry appears unjustified.
@@
    $\ell=5$. The refined triangle method [13] with $\ell_{1}=1$, $\ell_{2}=2$, $\ell_{3}=2$, leads to rectangular matrix multiplication $n\times m]\cdot[m\times m]$ or $n\times n^{2}]\cdot[n^{2}\times n^{2}]$ in the worst case. According to [39, Table 1], $\omega(1,2,2)=2\omega(1,0.5,1)\leq 2\cdot 2.043=4.086$, hence $T(n,m,5)=O(n^{4.086})$.
    On a side note, the extension method [13] with $j_{1}=3,j_{2}=2$, yields $T(n,m,5)=O(\alpha m\cdot m)=O(\alpha\cdot m^{2})$. Observe however, that $O(\alpha^{3}\cdot m)=O(\alpha\cdot m^{2})$, so the above upper bound does not beat that of the algorithm of Chiba and Nishizeki.
@@
    $\ell=7$. The refined triangle method [13] with $\ell_{1}=2$, $\ell_{2}=2$, $\ell_{3}=3$, leads to rectangular matrix multiplication $m\times m]\cdot[m\times\alpha m]$ or $m\times m]\cdot[m\times m^{3/2}]$ in the worst case. Since $\omega(1,1.5,1)\leq 2.795$, it follows that $T(n,m,7)=O(m^{2.795})$.
@@
    $\ell=8$. By [13, Thm. 2], it takes $O(m^{\beta(8)/2})$ time, where $\beta(8)=\omega(2,3,3)$. According to [1, Table 1], $\omega(1,0.6,1)\leq 2.0924$ and $\omega(1,0.7,1)\leq 2.1528$. By the convexity of $\omega(x,y,z)$ (see [25]), we have $\beta(8)=\omega(3,2,3)\leq\omega(1,0.6,1)+2\omega(1,0.7,1)\leq 6.398$. Thus, $T(n,m,8)=O(m^{\beta(8)/2})=O(m^{3.199})$.
@@
    $\ell=10$. By [13, Thm. 2], it takes $O(m^{\beta(10)/2})$ time, where $\beta(10)=\omega(3,3,4)$. Since according to [23, Table 3], $\omega(1,1.3,1)\leq 2.6217$ and $\omega(1,1.4,1)\leq 2.7084$, by the convexity of $\omega(x,y,z)$, we have $\omega(3,3,4)=\omega(3,4,3)=3\omega(1,1.3*2/3+1.4/3,1)\leq 2\omega(1,1.3,1)+\omega(1,1.4,1)\leq 7.9518$. Thus, $T(n,m,10)=O(m^{\beta(10)/2})=O(m^{3.976})$. Example: for graphs with $\alpha=O(m^{3/8})$, $K_{10}$’s can be counted/detected in $O(m^{3.636})$ time, by the last entry in the right column, whereas the other entries give $O(m^{3.976})$ time or more.
@@
    $\ell=11$. By [13, Thm. 2], it takes $O(m^{\beta(11)/2})$ time, where $\beta(11)=\omega(3,4,4)$. By [39, Table 1], we have $\omega(1,0.75,1)\leq 2.1863$, which yields $\omega(3,4,4)=4\omega(1,0.75,1)\leq 4\times 2.1863\leq 8.746$. Consequently, $T(n,m,11)=O(m^{\beta(11)/2})=O(m^{4.373})$.